OpenStudy (ksaimouli):

Venus rotates slowly about its axis, the period being 243 days. The mass of Venus is 4.87x10^24 kg. Determine the radius for a synchronous satellite in orbit about Venus.

OpenStudy (ksaimouli):

@JakeV8

OpenStudy (ksaimouli):

@asnaseer @cwrw238

OpenStudy (anonymous):

I used to be better at this :( Trying to resurrect some old knowledge...

OpenStudy (ksaimouli):

:-l

OpenStudy (anonymous):

You end up using the velocity of the satellite as v = d/t where d is the circumference of the orbit and t = 243 days (use appropriate units though). Then you can do something that I can't quite recall with radial velocity... v^2 / r = GM/r^2 I think is the formula, except that uses G from earth's gravity, so you would need to use the numbers for Venus instead. This might give you an idea though, but I realize it's not exactly a full explanation :)